OpenStudy (anonymous):

At a certain time during the day a tall tree casts a shadow of 32 feet. A shorter tree 12 feet tall casts a shadow of 16 feet. What is the height of the taller tree?

OpenStudy (anonymous):

A) 16 feet B) 20 feet C) 24 feet D) 28 feet

OpenStudy (anonymous):

12/16=x/32 x=12*32/16

OpenStudy (anonymous):

Thank you :) That makes sense.. You're amazing :D